y-4=-1/2(x-6) into standard form?
Answer by ankor@dixie-net.com(22740)

y-4 = -1/2(x-6) into standard form?
:
y - 4 = -.5x + 3
+.5x + y = 3 + 4
.5x + y = 7
Get rid of the decimal, mult by 10
5x + 10y = 70 is the standard form